OpenStudy (anonymous):

What is the square root of 18

OpenStudy (anonymous):

Google it man. Or use a calculator. You're wasting time by asking question like this one. Math.sqrt(18) = 4.242640687119285

OpenStudy (anonymous):

Maybe he asks for simplification? In that case; \(\sqrt{18}=\sqrt{9\cdot2}=\sqrt{9}\sqrt{2}=3\sqrt{2}\)

OpenStudy (anonymous):

Thank you